OpenStudy (anonymous):

What is the third term in the expansion of (x+2)^6?

OpenStudy (anonymous):

(x+2) (x+2) (x+2) (x+2) (x+2) (x+2) = (x^2 + 4x + 4) (x+2) (x+2) (x+2) (x+2) = (x^2 + 4x + 4) (x^2 + 4x + 4) (x +2) = can you do the rest?

OpenStudy (raden):

it can using : 6C2 (x)^(6-2) * 2^2